The effective mass of electron in the band is given by
1
Expert's answer
2021-10-05T10:06:05-0400

The effective mass of electron in the band is given by

"\frac{1}{m^*}=\frac{1}{\hbar^2}\frac{d^2E(k)}{dk^2}"